# Hilbert’s Theorem 90, periodicity, and roots of Artin-Schreier polynomials

### v2

@@ 4 Formulas for the roots of Artin-Schreier polynomials
Set $E={\mathbb{F}}_{2^{n}}$ and $F={\mathbb{F}}_{2}$. The polynomials in the last line of Table 4.7 are irreducible over $F$ (and hence can equal $m_{\zeta/F}(t)$), and $\zeta$ has order $12^{n_{2}}-$ 1. By Theorem 1.1, $x=R(y,z)$ is a root of $t^{2}+t+y$ where $z\in E$ is an element of degree $e=n_{2}$ satisfying $\textup{Tr}^{E}_{F}(z)=1$. Let $K$ be the subfield of $E$ with $|K:F|=n_{2}$ and $|E:K|=n_{2^{\prime}}$. We choose an element $\zeta\in K$ with $K={\mathbb{F}}_{2}(\zeta)$ and $\textup{Tr}^{K}_{F}(\zeta)=1$, and set $z=(n_{2^{\prime}})^{-1}\zeta$. Let $s_{i}=\sum_{j=0}^{i-1}\zeta^{2^{j}}$. Then $s_{0}=s_{2e}=0$ and $s_{i+2e}=s_{i}$, so it suffices to list $x_{i}=(n_{2^{\prime}})^{-1}s_{i}$ for $0\leqslant i<2e$. Indeed, since $s_{e}=1$ and $s_{i+e}=1+s_{i}$, it suffices to list $x_{i}$ for $0\leqslant i<e$.
(a) Suppose that $n_{2}=1$. Either set $\zeta=1$, $z=\frac{1}{n}$, or set $p=2$ in Corollary 4.2. Then $x^{2}+x+y=0$ where $x\kern-1.0pt=\kern-1.0pt\sum_{i=0}^{n-1}\frac{i}{n}y^{2^{i}}\kern-1.0pt=\kern-1.0pt\sum_{j=0}^{(n-1)/2}y^{2^{2j+1}}$ and $x_{i}=0,1$ for $i\equiv 0,1\pmod{2}$.
(b) Suppose that $n_{2}=2$. Set $\zeta=\omega$ or set $p=2$ into Corollary 4.6. Since $2\nmid(n/2)$ the value of $x_{i}=(\frac{n}{2})^{-1}\left(\lfloor\frac{i}{2}\rfloor-r_{i}\omega\right)$ is $0,\omega,1,\omega^{2}$ when $i\equiv 0,1,2,3\pmod{4}$. (Recall that $r_{i}=i-2\lfloor\frac{i}{2}\rfloor\in\{0,1\}$.) Equivalent formulas for $x$ in cases (a) and (b) were given previously by Chen [1, Theorems 1, 2] and by Lahtonen [4].
(c) Suppose that $n_{2}=4$. Then $\zeta=\alpha\in{\mathbb{F}}_{16}$ has order $12^{4}-$ 1 and $\textup{Tr}^{K}_{F}(\zeta)=1$. Further, the values of $s_{i}=\sum_{j=0}^{i-1}\zeta^{2^{i}}$ are $0,\alpha,\alpha^{13},\alpha^{6},1,\alpha^{12},\alpha^{7},\alpha^{8}$ for $i\equiv 0,1,\dots,7\pmod{8}$.
(d) Suppose that $n_{2}=8$. Then $\zeta=\beta\in{\mathbb{F}}_{256}$ has order $12^{8}-$ 1 and satisfies $\textup{Tr}^{K}_{F}(\zeta)=1$. Furthermore, the values of $s_{i}=\sum_{j=0}^{i-1}\zeta^{2^{i}}$ are $0,\beta,\beta^{100},\beta^{189},\beta^{29},\beta^{60},\beta^{154},\beta^{177}$ for $i\equiv 0,1,\dots,7\pmod{16}$. Also $s_{i+8}=1+s_{i}$ for $i\equiv 0,1,\dots,7\pmod{16}$.
(e) Suppose that $n_{2}=16$. Then $\zeta=\gamma\in{\mathbb{F}}_{2^{16}}$ has order $12^{16}-$ 1 and $\textup{Tr}^{K}_{F}(\zeta)=1$. In addition, $s_{i}=\sum_{j=0}^{i-1}\zeta^{2^{j}}$ equals $0,\gamma,\gamma^{64409},\gamma^{48754},\gamma^{27742},\gamma^{48469},\gamma^{1146},\gamma^{22404},\gamma^{64313},\gamma^{47682},$ $\gamma^{63219},\gamma^{45929},\gamma^{55680},\gamma^{46875},\gamma^{7495},\gamma^{32204}$ and $s_{i+16}=1+s_{i}$ for $i\equiv 0,1,\dots,15\pmod{32}$. ∎
@@
If $e=n_{p}$ and $F={\mathbb{F}}_{p}$, then there are $p^{e-1}$ choices for an element $\zeta\in K={\mathbb{F}}_{p^{e}}$ with $\textup{Tr}^{K}_{F}(\zeta)=1$. In Lemma 4.7 we chose $\zeta$ so that (a) $\sum_{i=0}^{n_{2}-1}\zeta^{2^{i}}=1$, (b) the minimal polynomial $m_{\zeta/F}(t)$ has many zero coefficients, and (c) $\zeta$ has multiplicative order $12^{e}-$ 1. There are many choices for $\zeta$ with these three properties, and there appears to be no canonical choice in general.

### v3

@@ 4 Formulas for the roots of Artin-Schreier polynomials
Set $E={\mathbb{F}}_{2^{n}}$, $F={\mathbb{F}}_{2}$ and $e=n_{2}$. The last line of Table 1 lists big irreducible polynomials over $F$ whose roots have order $2^{n_{2}}-1$, see Lemma 4.6 and Remark 4.10. The polynomials are big by Definition 4.5. The irreducibility over $F$, and the order of $z$, are easily checked using a computational algebra system. These together imply that $K\coloneq F(z)$ has $|K:F|=n_{2}$ and $\textup{Tr}^{K}_{F}(z)=1$. Hence $\textup{Tr}^{E}_{F}(z)=|E:K|=n_{2^{\prime}}=1$ as desired. By Theorem 1.1, $x=R(y,z)$ is a root of $t^{2}+t+y$. In order to use our formulas, the reader will need to know the values of $x_{i}=\sum_{j=0}^{i-1}z^{2^{j}}$. Since $x_{0}=x_{2e}=0$ and $x_{i+2e}=x_{i}$, it suffices to list $x_{i}$ for $0\leqslant i<2e$. However, since $x_{e}=\textup{Tr}_{F}^{K}z=1$ and $x_{i+e}=1+x_{i}$, it suffices to list $x_{i}$ for $0\leqslant i<e$.
(a) Suppose that $n_{2}=1$. Set $q=p=2$ in Corollary 4.2. Then $x^{2}+x+y=0$ where $x=\sum_{i=0}^{n-1}\frac{i}{n}y^{2^{i}}=\sum_{j=0}^{(n-1)/2}y^{2^{2j+1}}$. Here $x_{i}\in\{0,1\}$ satisfies $x_{i}\equiv i\pmod{2}$.
(b) Suppose that $n_{2}=2$. Set $p=2$ and $z=\omega$ into Lemma 4.8. Since $n/2$ is odd the value of $x_{i}=(\frac{n}{2})^{-1}\left(\lfloor\frac{i}{2}\rfloor-r_{i}\omega\right)$ is $0,\omega,1,\omega^{2}$ when $i\equiv 0,1,2,3\pmod{4}$; recall that $r_{i}=i-2\lfloor\frac{i}{2}\rfloor\in\{0,1\}$. Equivalent formulas for $x$ in cases (a) and (b) were given previously by Chen [1, Theorems 1, 2] and by Lahtonen [5].
(c) Suppose that $n_{2}=4$. Then $z=\alpha\in{\mathbb{F}}_{16}$ has order $2^{4}-1$ and $\textup{Tr}^{K}_{F}(z)=1$. Also the values of $x_{i}=\sum_{j=0}^{i-1}z^{2^{i}}$ are $0,\alpha,\alpha^{13},\alpha^{6},1,\alpha^{12},\alpha^{7},\alpha^{8}$ for $i\equiv 0,1,\dots,7\pmod{8}$.
(d) Suppose that $n_{2}=8$. Then $z=\beta\in{\mathbb{F}}_{256}$ has order $2^{8}-1$ and $z$ satisfies $\textup{Tr}^{K}_{F}(z)=1$. Furthermore, the values of $x_{i}=\sum_{j=0}^{i-1}z^{2^{i}}$ are $0,\beta,\beta^{100},\beta^{189},\beta^{29},\beta^{60},\beta^{154},\beta^{177}$ for $i\equiv 0,1,\dots,7\pmod{16}$. Also $x_{i+8}=1+x_{i}$ for $i\equiv 0,1,\dots,7\pmod{16}$.
(e) Suppose that $n_{2}=16$. Then $z=\gamma\in{\mathbb{F}}_{2^{16}}$ has order $2^{16}-1$ and $\textup{Tr}^{K}_{F}(z)=1$. In addition, $x_{i}=\sum_{j=0}^{i-1}z^{2^{j}}$ equals $0,\gamma,\gamma^{64409},\gamma^{48754},\gamma^{27742},\gamma^{48469},\gamma^{1146},\gamma^{22404},\gamma^{64313},\gamma^{47682},$ $\gamma^{63219},\gamma^{45929},\gamma^{55680},\gamma^{46875},\gamma^{7495},\gamma^{32204}$ and $x_{i+16}=1+x_{i}$ for $i\equiv 0,1,\dots,15\pmod{32}$.
(f) Suppose that $n_{2}=32$. Then $z=\delta\in{\mathbb{F}}_{2^{32}}$ has order $2^{32}-1$ and $\textup{Tr}^{K}_{F}(z)=1$. Writing $x_{i}=\sum_{j=0}^{i-1}z^{2^{j}}$ as a power of $\delta$ for $i\equiv 0,1,\dots,15\pmod{32}$, involves solving the discrete logarithm problem. This was done in the previous cases as writing a power of a primitive element was shorter than writing polynomials in the primitive element. ∎
@@
It is an open question whether big primitive polynomials of degree $e=n_{2}=2^{k}$ in the last line of Table 1 exist for arbitrarily large choices of $k$. This turns out to be true if $2^{e}-1$ is square-free because by Lemma 4.11 the cyclotomic polynomial $\Phi_{2^{e}-1}(t)$ is big. This follows because $\Phi_{2^{e}-1}(t)$ is a tensor product [2] $\Phi_{p_{1}}(t)\otimes\cdots\otimes\Phi_{p_{r}}(t)$ of big polynomials where $2^{e}-1=p_{1}\cdots p_{r}$ and the $p_{i}$ are distinct primes. However, it is not known whether $2^{2^{k}}-1=\prod_{i=0}^{k-1}(2^{2^{i}}+1)$ is square-free for all $k\geqslant 1$. (The Fermat numbers $2^{2^{i}}+1$ are pairwise coprime, but it not known whether Fermat numbers are square-free.) The requirement that the polynomial be primitive (i.e. a root $z$ has order $2^{e}-1$) is not necessary, we assumed primitivity only so that the non-zero elements $x_{i}$ could be written as a power of $z$. What surprised the author is that a root of the Schreier polynomial $t^{2}+t+y$ in ${\mathbb{F}}_{2^{n}}$ can always be described using only ${\rm O}(e^{2})$ bits of information.
Lemma 4.11.
The tensor product of big nonzero polynomials is big, see Definition 4.5.
Proof.
Let $a(t)$ and $b(t)$ be big nonzero polynomials of respective degrees $m$ and $n$. The result is clear if $mn=0$ by [2, p. 307]. Suppose that $mn>0$ and $a(t)=\prod_{i=1}^{m}(t-\alpha_{i})$ and $b(t)=\prod_{j=1}^{n}(t-\beta_{j})$ are monic with roots $\alpha_{i},\beta_{j}$ in an extension field. By Definition 4.5, $a(t)=t^{m}+a_{m-1}t^{m-1}+\cdots$ and $b(t)=t^{n}+b_{n-1}t^{n-1}+\cdots$ where $a_{m-1}=-\sum_{i=1}^{m}\alpha_{i}\neq 0$ and $b_{n-1}=-\sum_{j=1}^{n}\beta_{j}\neq 0$. However, $(a\otimes b)(t)=\prod_{i,j}(t-\alpha_{i}\beta_{j})$ by [2, p. 307], and hence $(a\otimes b)(t)=t^{mn}-(\sum_{i,j}\alpha_{i}\beta_{j})t^{mn-1}+\cdots$. Therefore $(a\otimes b)(t)$ is big since $\sum_{i,j}\alpha_{i}\beta_{j}=a_{m-1}b_{n-1}\neq 0$. More generally, if $a(t)$ and $b(t)$ have leading terms $a_{m}$ and $b_{n}$ respectively, then $(a\otimes b)(t)$ has leading term $a_{m}^{n}b_{n}^{m}\neq 0$ by [2, p. 307], and hence is big. ∎